OpenStudy (anonymous):

exponetial form x^6/2

OpenStudy (anonymous):

\[x ^{6/2}\]

OpenStudy (anonymous):

x^3

OpenStudy (anonymous):

choices are\[\sqrt[6]{x ^{2}}...\sqrt{x ^{3}}...x ^{3}....\sqrt[3]{x}\]